Operations Report – January 07th

Crew 219 Operations Report 07-JAN-2020

Sol: 2

Name of person filing report: Alejandro Luis Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 13.5

From what time last night: 17:00

To what time this morning: 06:30

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 49%

Diesel Reading – 74%

Station Propane Reading- 52%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147447,5

Water (static tank) (gal) – 526

Static to loft Pump used – Yes

Water in Green Hab (gal): 204.88

Water in Science Dome (gal): 0

Toilet Tank Emptied – No

Diemos rover used: Still in the Work shop


Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Assigned to Director


Beginning Charge:

End Charge:

Currently Charging:

Spirit Rover used- Yes

EVA #4

Starting Hours: 123.3

Beginning Charge: 100%

Ending Hours: 123.5

Ending Charge: 67%

Opportunity Rover used: Still in workshop


Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes

EVA #4

Starting Hours: 129.1

Beginning Charge: 100%

Ending Hours: 129.3

Ending Charge: 71 %

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where? Yes, used to fill static tank by Outpost.

CrewCar used and why, where? No

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Suits #1 and #7 are fully functional, weak air flow from yesterday’s EVA could have been from the low temperatures. Suit #5 does not work. The multimeter was used to see where power is not being transmitted and found that all connection from battery to circuit board are good. However, there is no power being delivered to the fans. Suit #3 has weak air flow coming from both fans even on full power. Battery testing on Suit #3 will be conducted tomorrow. Red Helmet was fixed by securing the head pad back in place using cable ties. The connector has been placed on the helmet missing a connector, however, there is no gasket for this connector. Also, one of the chargers got disconnected from its connector. This was fixed by soldering the cable back to the connector and this charger is working again.

Summary of Hab: Heating strip is working to keep the pipe from the static tank to loft tank from freezing. However, the insulation looks to be falling apart from the cold temperatures. Also, the cable obstructing the fire escape window has been removed and placed inside the RAM.

Summary of Science Dome operations: Nominal

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: Requesting Teflon tape for the loft tank feedline, new insulation for the Static tank to Loft tank pipe, and gasket for one of the 2 piece helmet connectors.

Copyright © The Mars Society. All rights reserved. | Main Site